the resultant vector P AND Q IS R IF Q IS DOUBLED R IS DOUBLED WHEN Q IS REVERSED R IS AGAIN DOUBLED FIND P:Q:R

Grade:Select Grade

1 Answers

Arun
6 years ago
Dear student
p^2 + q^2 + 2pq cos @ = r^2           ...(i)
p^2 + 4q^2 + 4pq cos @ = 4r^2           …...(ii)
p^2 + q^2 – 2pq cos @ = 4r^2        …....(iii)
equation (i) – (iii) 
pq cos@ =  –3 r^2 /4
p^2 + q^2 = 5/2 * r^2
p^2 + 4 q^2 = 7 r^2
p^2 + q^2 = 5/2 * r^2
on solving
p^2 = r^2
and
q^2 = 3/2 * r^2
hence
p^2 : q^2: r^2 = 2 : 3: 2
p : q : r  = \small \sqrt2 : \small \sqrt3 : \small \sqrt2
Regards
Arun (askIITians forum expert)